# Fast and Robust Visuomotor Riemannian Flow Matching Policy

### v1

@@ Abstract
Diffusion-based visuomotor policies excel at learning complex robotic tasks by effectively combining visual data with high-dimensional, multi-modal action distributions. However, diffusion models often suffer from slow inference due to costly denoising processes or require complex sequential training arising from recent distilling approaches. This paper introduces Riemannian Flow Matching Policy (RFMP), a model that inherits the easy training and fast inference capabilities of flow matching (FM). Moreover, RFMP inherently incorporates geometric constraints commonly found in realistic robotic applications, as the robot state resides on a Riemannian manifold. To enhance the robustness of RFMP, we propose Stable RFMP (SRFMP), which leverages LaSalle’s invariance principle to equip the dynamics of FM with stability to the support of a target Riemannian distribution. Rigorous evaluation on eight simulated and real-world tasks show that RFMP successfully learns and synthesizes complex sensorimotor policies on Euclidean and Riemannian spaces with efficient training and inference phases, outperforming Diffusion Policies while remaining competitive with Consistency Policies.
@@ I Introduction
In this paper, we demonstrate the effectiveness of RFMP to learn complex real-world visuomotor policies and present a systematic evaluation of the performance of RFMP on both simulated and real-world manipulation tasks. Moreover, we propose Stable Riemannian Flow Matching Policy (SRFMP) to enhance the robustness of RFMP (see Figure 1). SRFMP builds on stable flow matching (SFM) [19, 20], which leverages LaSalle’s invariance principle [21] to equip the dynamics of FM with stability to the support of the target distribution. Unlike SFM, which is limited to Euclidean spaces, SRFMP generalizes this concept to Riemannian manifolds, guaranteeing the stability of the RFM dynamics to the support of a Riemannian target distribution. We systematically evaluate RFMP and SRFMP across $8$ tasks in both simulation and real-world settings, with policies conditioned on both state and visual observations. Our experiments demonstrate that RFMP and SRFMP inherit the advantages from FM models, achieving comparable performance to DP with fewer evaluation steps (i.e., faster inference) and significantly shorter training times. Moreover, our models achieve competitive performance compared to CP [3] on several simulated tasks. Notably, SRFMP requires fewer ODE steps than RFMP to achieve an equivalent performance, resulting in even faster inference times.
In summary, beyond demonstrating the effectiveness of our early work on simulated and real robotic tasks, the main contributions of this article are threefold: (1) We introduce Stable Riemannian Flow Matching (SRFM) as an extension of SFM [20] to incorporate stability into RFM; (2) We propose stable Riemannian flow matching policy (SRFMP), which combines the easy training and fast inference of RFMP with stability guarantees to Riemannian target action distribution; (3) We systematically evaluate both RFMP and SRFMP across $8$ tasks from simulated benchmarks and real settings. Supplementary material is available on the paper website https://sites.google.com/view/rfmp.
@@ II Related work
Diffusion Models [28] recently became state-of-art in imitation learning because of their ability to learn multi-modal and high-dimensional action distributions. They have been primarily employed to learn motion planners [29], and complex control policies [2, 30, 4]. Recent extensions consider to use 3D visual representations from sparse point clouds [6], and to employ equivariant networks for learning policies that, by design, are invariant to changes in scale, rotation, and translation [7]. However, a major drawback of diffusion models is their slow inference process. In [2], DP requires $10$ to $100$ denoising steps, i.e., $0.1$ to $1$ second on a standard GPU, to generate each action. Consistency models (CM) [31] arise as a potential solution to overcome this drawback [3, 32, 33]. CM distills a student model from a pretrained diffusion model (i.e., a teacher), enabling faster inference by establishing direct connections between points along the probability path. Nevertheless, this sequential training process increases the overall complexity and time of the whole training phase.
@@ IV-A Riemannian Flow Matching Policy
Algorithm 1 RFMP Training Process
@@ IV-B1 Stable Euclidean FMP
$u(\bm{\xi}_{t}|\bm{\xi}_{1})=\left[\begin{aligned} u_{{\bm{x}}}({\bm{x}}_{t}|{\bm{x}}_{1})\\ u_{\tau}(\tau_{t}|\tau_{1})\end{aligned}\right]=\left[\begin{aligned} -\lambda_{\bm{x}}({\bm{x}}_{0}-{\bm{x}}_{1})\\ -\lambda_{\tau}(\tau_{0}-\tau_{1})\end{aligned}\right],$ (16)
@@
Figure 2: Flows of the RFM (top) and SRFM (bottom) trained on the $\mathsf{L}$-shape LASA dataset projected on the sphere manifold. Orange points represent the training dataset, while blue points show generated probability path at different times $t=\{0.0,1.0,1.5\}$ across the three columns.
@@ 1 Training Process
11 Sample ${\bm{a}}_{0}\sim p_{0}$, and set $t=0$, $\bm{\xi}_{0}=[{\bm{a}}_{0},\tau_{0}]$.
12 while $t\leq N$ do
13       Update time step $t=t+1$,
14       if $t=1$ then
@@ 15             $\Delta t=\frac{1}{\lambda_{{\bm{x}}}}$
17             $\Delta t=\varepsilon<\frac{1}{\lambda_{{\bm{x}}}}$
18      Integration on the learned Riemannian vector field $\bm{\xi}_{t}=\text{Exp}_{\bm{\xi}_{t-1}}(v_{{\bm{\theta}}}(\bm{\xi}_{t-1},{\bm{o}})\Delta t)$
@@ 17             $\Delta t=\varepsilon<\frac{1}{\lambda_{{\bm{x}}}}$
$u(\bm{\xi}_{t}|\bm{\xi}_{1})=\left[\begin{aligned} u_{\bm{x}}({\bm{x}}_{t}|{\bm{x}}_{1})\\ u_{\tau}(\tau_{t}|\tau_{1})\end{aligned}\right]=\left[\begin{aligned} -\lambda_{\bm{x}}\text{Log}_{{\bm{x}}_{1}}({\bm{x}}_{0})\\ -\lambda_{\tau}(\tau_{0}-\tau_{1})\end{aligned}\right],$ (20)
@@
The parameters $\lambda_{\bm{x}}$ and $\lambda_{\tau}$ have the same influence as in the Euclidean case. Notice that the spatial part of the stable flow (21) closely resembles the geodesic flow (8) proposed in [16]. Figure 2 shows an example of learned RFM and SRFM flows at times $t=\{0,1,1.5\}$. The RFM flow diverges from the target distribution at times $t>1$, while the SRFM flow is stable and adheres to the target distribution for $t\geq 1$.
@@
This approach, hereinafter referred to as stable Riemannian flow matching policy (SRFMP), is summarized in Algorithm 2. The learned SRFMP vector field drives the flow to converge to the target distribution within a certain time horizon, ensuring that it remains within this distribution, as illustrated by the bottom row of Figures 1 and 2. In contrast, the RFMP vector field may drift the flow away from the target distribution at $t>1$ (see Figures 1 and 2, top row). Therefore, SRFMPs provide flexibility and increased robustness in designing the generation process, while RFMPs are more sensitive to the integration process.
@@ V Experiments
We thoroughly evaluate the performance of RFMP and SRFMP on a set of six simulation settings and two real-world tasks. The simulated benchmarks are: (1) The Push-T task from [2]; (2) A Sphere Push-T task, which we introduce as a Riemannian benchmark; and (3)-(6) Four tasks (Lift, Can, Square, and Tool Hang) from the large-scale robot manipulation benchmark Robomimic [10]. The real-world robot tasks correspond to: (1) A Pick & Place task; and (2) A Mug Flipping task. Collectively, these eight tasks serve as a benchmark to evaluate (1) the performance, (2) the training time, and (3) the inference time of RFMP and SRFMP with respect to state-of-the-art generative policies, i.e., DP and extensions thereof.
@@ V-A1 RFMP and SRFMP Implementation
Our RFMP implementation builds on the RFM framework from Chen and Lipman [16]. We parameterized the vector field $v_{t}(\bm{a}|\bm{o};\bm{\theta})$ using the UNet architecture employed in DP [2], which consists of $3$ layers with downsampling dimensions of $(256,512,1024)$ and $(128,256,512)$ for simulated and real-world tasks. Each layer employs a $1$-dimensional convolutional residual network as proposed in [29]. We implement a Feature-wise Linear Modulation (FiLM) [58] to incorporate the observation condition vector ${\bm{o}}$ and time step $t$ into the UNet. Instead of directly feeding the FM time step $t$ as a conditional variable, we first project it into a higher-dimensional space using a sinusoidal embedding module, similarly to DP. For tasks with image-based observations, we leverage the same vision perception backbone as in DP [2]. Namely, we use a standard ResNet-$18$ in which we replace: (1) the global average pooling with a spatial softmax pooling, and (2) BatchNorm with GroupNorm. Our SRFMP implementation builds on the SFM framework [20]. We implement the same UNet as RFMP to represent $v_{{\bm{x}}}$ by replacing the time step $t$ by the temperature parameter $\tau$. We introduce an additional Multi-Layer Perceptron (MLP) to learn $v_{\tau}$. As for $t$ in RFMP, we employed a sinusoidal embedding for the input $\tau$. The boundaries $\tau_{0}$ and $\tau_{1}$ are set to 00 and $1$ in all experiments.
We implement different prior distributions for different tasks. For the Euclidean Push-T and the four Robomimic tasks, the action space is $\mathbb{R}^{d}$ and we thus define the prior distribution as a Euclidean Gaussian distribution for both RFMP and SRFMP. For the Sphere Push-T, the action space is the hypersphere $\mathcal{S}^{2}$. In this case, we test two types of Riemannian prior distribution, namely a spherical uniform distribution, and a wrapped Gaussian distribution [59, 60], illustrated in Figure 3. Regarding the real robot tasks, the task space is defined as the product manifold $\mathcal{M}=\mathbb{R}^{3}\times\mathcal{S}^{3}\times\mathbb{R}^{1}$, whose components represent the position, orientation (encoded as quaternions), and opening of the gripper. The Euclidean and hypersphere parts employ Euclidean Gaussian distributions and a wrapped Gaussian distribution, respectively.
@@
State-based Robomimic
@@
Vision-based Robomimic
@@
For all experiments, we optimize the network parameters of RFMP and SRFMP using AdamW [61] with a learning rate of $\eta=1\times 10^{-4}$ and weight decay of $w_{d}=0.001$ based on an exponential moving averaging (EMA) framework on the weights [62] with a decay of $w_{\text{EMA}}=0.999$. Moreover, we use an action prediction horizon $T_{p}=16$, an action horizon $T_{a}=T_{p}/2$, and an observation horizon $T_{o}=2$. We set the SRFMP parameters as $\lambda_{{\bm{x}}}=\lambda_{\tau}=2.5$. Table I summarizes the image resolution, number of parameters, and number of training epochs used in each experiment.
@@ V-A2 Baselines
In the original DP paper [2], the policy is trained using either Denoising Diffusion Probabilistic Model (DDPM) [12] or Denoising Diffusion Implicit Model (DDIM) [13]. In this paper, we prioritize faster inference and thus employ DDIM-based DP for all our experiments. We train DDIM with $100$ denoising steps. The prior distribution is a standard Gaussian distribution unless explicitly mentioned. During training we use the same noise scheduler as in [2], the optimizer AdamW with the same learning rate and weight decay as for RFMP and SRFMP. Note that DP does not handle data on Riemannian manifolds, and thus does not guarantee that the resulting trajectories lie on the manifold of interest for tasks with Riemannian action spaces, e.g., the Sphere Push-T and real-world robot experiments. In these cases, we post-process the trajectories obtained during inference and project them on the manifold. In the case of the hypersphere manifold, the projection corresponds to a unit-norm normalization. We also compare RFMP and SRFMP against CP [3] on the Robomimic tasks with vision-based observations. To do so, we use the performance values reported in [3].

### v2

@@ Abstract
Diffusion-based visuomotor policies excel at learning complex robotic tasks by effectively combining visual data with high-dimensional, multi-modal action distributions. However, diffusion models often suffer from slow inference due to costly denoising processes or require complex sequential training arising from recent distilling approaches. This paper introduces Riemannian Flow Matching Policy (RFMP), a model that inherits the easy training and fast inference capabilities of flow matching (FM). Moreover, RFMP inherently incorporates geometric constraints commonly found in realistic robotic applications, as the robot state resides on a Riemannian manifold. To enhance the robustness of RFMP, we propose Stable RFMP (SRFMP), which leverages LaSalle’s invariance principle to equip the dynamics of FM with stability to the support of a target Riemannian distribution. Rigorous evaluation on eight simulated and real-world tasks show that RFMP successfully learns and synthesizes complex sensorimotor policies on Euclidean and Riemannian spaces with efficient training and inference phases, outperforming Diffusion Policies and Consistency Policies.
@@ I Introduction
In this paper, we demonstrate the effectiveness of RFMP to learn complex real-world visuomotor policies and present a systematic evaluation of the performance of RFMP on both simulated and real-world manipulation tasks. Moreover, we propose Stable Riemannian Flow Matching Policy (SRFMP) to enhance the robustness of RFMP, as RFMP performance can be sensitive to the ODE integration horizon during inference (see Figures 1 and 2). SRFMP builds on stable flow matching (SFM) [19, 20], which leverages LaSalle’s invariance principle [21] to equip the dynamics of FM with stability to the support of the target distribution. Unlike SFM, which is limited to Euclidean spaces, SRFMP generalizes this concept to Riemannian manifolds, guaranteeing the stability of the RFM dynamics to the support of a Riemannian target distribution. We systematically evaluate RFMP and SRFMP across $10$ tasks in both simulation and real-world settings, with policies conditioned on both state and visual observations. Our experiments demonstrate that RFMP and SRFMP inherit the advantages from FM models, achieving comparable performance to DP with fewer evaluation steps (i.e., faster inference) and significantly shorter training times. Moreover, under the same training epochs, our methods outperform both DP and CP. Notably, SRFMP requires fewer ODE steps than RFMP to achieve an equivalent performance, resulting in even faster inference times.
In summary, beyond demonstrating the effectiveness of our early work on simulated and real robotic tasks, the main contributions of this article are threefold: (1) We introduce Stable Riemannian Flow Matching (SRFM) as an extension of SFM [20] to incorporate stability into RFM; (2) We propose stable Riemannian flow matching policy (SRFMP), which combines the easy training and fast inference of RFMP with stability guarantees to a Riemannian target action distribution; (3) We systematically evaluate both RFMP and SRFMP across $10$ tasks from simulated benchmarks and real settings. Supplementary material is available on the paper website https://sites.google.com/view/rfmp.
@@ II Related work
Diffusion Models [28] recently became state-of-art in imitation learning because of their ability to learn multi-modal and high-dimensional action distributions. They have been primarily employed to learn motion planners [29], and complex control policies [2, 30, 4]. Recent extensions use 3D visual representations from sparse point clouds [6], and employ equivariant networks for learning policies that, by design, are invariant to changes in scale, rotation, and translation [7]. However, a major drawback of diffusion models is their slow inference process. In [2], DP requires $10$ to $100$ denoising steps, i.e., $0.1$ to $1$ second on a standard GPU, to generate each action. Consistency models (CM) [31] arise as a potential solution to overcome this drawback [3, 32, 33]. CM distills a student model from a pretrained diffusion model (i.e., a teacher), enabling faster inference by establishing direct connections between points along the probability path. Nevertheless, this sequential training process increases the overall complexity and time of the whole training phase.
@@ 1 Training Process
9Inference Step
10 Sample ${\bm{a}}_{0}\sim p_{0}$, and set $t=0$.
11 while $t\leq 1$ do
12       Integration on the learned Riemannian vector field $\bm{\xi}_{t+\Delta t}=\text{Exp}_{\bm{\xi}_{t}}(v_{{\bm{\theta}}}(\bm{\xi}_{t},{\bm{o}})\Delta t)$
13       Update time $t=t+\Delta t$
14
Algorithm 1 RFMP Training & Inference Process
@@ 13       Update time $t=t+\Delta t$
Although RFM is theoretically designed to match a time-dependent vector field defined over the time horizon $t\in[0,1]$, we observed that, in practice, its inference performance often improves when evaluating trajectories slightly beyond $t=1$. To investigate this phenomenon, we conduct experiments on two tasks: The Euclidean Push-T task from [2] and the Square task from the Robomimic robotic manipulation benchmark [10]. As shown in Figure 2, RFMP performance improves when the ODE integration horizon slightly exceeds $t=1$, but it worsens and eventually gets unstable when integrating for longer time horizons. This observation motivates the development of SRFMP introduced next, which enhances RFMP by explicitly enforcing inference stability beyond the unit interval using the LaSalle’s invariance principle. This approach stabilizes the policy inference, making it robust to the ODE integration horizon, as illustrated in Figures 13 and 4.
Figure 2: RFMP performance when extending inference time beyond $t=1$ on Euclidean PushT and Robomimic Square tasks.
@@ IV-B Stable Riemannian Flow Matching Policy
Figure 3: Flows of the RFM (top) and SRFM (bottom) trained on the $\mathsf{L}$-shape LASA dataset projected on the sphere manifold. Orange points represent the training dataset, while blue points are sampled from the generated probability path at different times $t=\{0.0,1.0,1.5\}$ across the three columns.
@@ IV-B1 Stable Euclidean FMP
$u(\bm{\xi}_{t}|\bm{\xi}_{1})=\left[\begin{aligned} u_{{\bm{x}}}({\bm{x}}_{t}|{\bm{x}}_{1})\\ u_{\tau}(\tau_{t}|\tau_{1})\end{aligned}\right]=\left[\begin{aligned} -\lambda_{\bm{x}}({\bm{x}}_{t}-{\bm{x}}_{1})\\ -\lambda_{\tau}(\tau_{t}-\tau_{1})\end{aligned}\right],$ (16)
@@
Figure 4: Flows of the RFM (top) and SRFM (bottom) on the SPD manifold $\mathcal{S}_{++}^{2}$. Orange points represent the training dataset, while blue points correspond to sampled from the generated probability path at different times $t=\{0.0,1.0,1.5\}$ across the three columns.
@@ 1 Training Process
11 Sample ${\bm{a}}_{0}\sim p_{0}$, and set $k=1$, $t=0$, $\bm{\xi}_{0}=[{\bm{a}}_{0},\tau_{0}]$.
12 while $k\leq N$ do
13
14      if $k=1$ then
@@ 15             $\Delta t=\frac{1}{\lambda_{{\bm{x}}}}$
17             $\Delta t=\varepsilon\leq\frac{1}{\lambda_{{\bm{x}}}}$
18      Integration on the learned Riemannian vector field $\bm{\xi}_{t+\Delta t}=\text{Exp}_{\bm{\xi}_{t}}(v_{{\bm{\theta}}}(\bm{\xi}_{t},{\bm{o}})\Delta t)$,
19       Update time $t=t+\Delta t$
20       Update iteration $k=k+1$
@@
$u(\bm{\xi}_{t}|\bm{\xi}_{1})=\left[\begin{aligned} u_{\bm{x}}({\bm{x}}_{t}|{\bm{x}}_{1})\\ u_{\tau}(\tau_{t}|\tau_{1})\end{aligned}\right]=\left[\begin{aligned} -\lambda_{\bm{x}}\text{Log}_{{\bm{x}}_{1}}({\bm{x}}_{t})\\ -\lambda_{\tau}(\tau_{t}-\tau_{1})\end{aligned}\right],$ (20)
@@
The parameters $\lambda_{\bm{x}}$ and $\lambda_{\tau}$ have the same influence as in the Euclidean case. Notice that the spatial part of the stable flow (21) closely resembles the geodesic flow (8) proposed in [16]. Figures 3 and 4 show examples of learned RFM and SRFM flows at times $t=\{0,1,1.5\}$ on the sphere and symmetric positive-definite (SPD) matrices manifold. The RFM flow diverges from the target distribution at times $t>1$, while the SRFM flow is stable and adheres to the target distribution for $t\geq 1$.
@@
This approach, hereinafter referred to as stable Riemannian flow matching policy (SRFMP), is summarized in Algorithm 2. The learned SRFMP vector field drives the flow to converge to the target distribution within a certain time horizon, ensuring that it remains within this distribution, as illustrated by the bottom row of Figures 13, and 4 . In contrast, the RFMP vector field may drift the flow away from the target distribution at $t>1$ (see the top row of Figures 13, and 4). Therefore, SRFMPs provide flexibility and increased robustness in designing the generation process, while RFMPs are more sensitive to the integration process.
@@ V Experiments
We thoroughly evaluate the performance of RFMP and SRFMP on a set of eight simulation settings and two real-world tasks. The simulated benchmarks are: (1) The Push-T task from [2]; (2) A Sphere Push-T task, which we introduce as a Riemannian benchmark; (3)-(7) Five tasks (Lift, Can, Square, Tool Hang, and Transport) from the large-scale robot manipulation benchmark Robomimic [10]; and (8) the Franka Kitchen benchmark [58] featuring complex, long-horizon tasks. The real-world robot tasks correspond to: (1) A Pick & Place task; and (2) A Mug Flipping task. Collectively, these ten tasks serve as a benchmark to evaluate (1) the performance, (2) the training time, and (3) the inference time of RFMP and SRFMP with respect to state-of-the-art generative policies, i.e., DP and extensions thereof.
@@ V-A1 RFMP and SRFMP Implementation
Our RFMP implementation builds on the RFM framework from Chen and Lipman [16]. We parameterize the vector field $v_{t}(\bm{a}|\bm{o};\bm{\theta})$ using the UNet architecture employed in DP [2], which consists of $3$ layers with downsampling dimensions of $(256,512,1024)$ and $(128,256,512)$ for simulated and real-world tasks. Each layer employs a $1$-dimensional convolutional residual network as proposed in [29]. We implement a Feature-wise Linear Modulation (FiLM) [59] to incorporate the observation condition vector ${\bm{o}}$ and time step $t$ into the UNet. Instead of directly feeding the FM time step $t$ as a conditional variable, we first project it into a higher-dimensional space using a sinusoidal embedding module, similarly to DP. For tasks with image-based observations, we leverage the same vision perception backbone as in DP [2]. Namely, we use a standard ResNet-$18$ in which we replace: (1) the global average pooling with a spatial softmax pooling, and (2) BatchNorm with GroupNorm. Our SRFMP implementation builds on the SFM framework [20]. We implement the same UNet as RFMP to represent $v_{{\bm{x}}}$ by replacing the time step $t$ by the temperature parameter $\tau$. We introduce an additional Multi-Layer Perceptron (MLP) to learn $v_{\tau}$. As for $t$ in RFMP, we employed a sinusoidal embedding for the input $\tau$. The boundaries $\tau_{0}$ and $\tau_{1}$ are set to 00 and $1$ in all experiments.
We implement different prior distributions for different tasks. For the Euclidean Push-T and the Robomimic tasks, the action space is $\mathbb{R}^{d}$ and we thus define the prior distribution as a Euclidean Gaussian distribution for both RFMP and SRFMP. For the Sphere Push-T, the action space is the hypersphere $\mathcal{S}^{2}$. In this case, we test two types of Riemannian prior distribution, namely a spherical uniform distribution, and a wrapped Gaussian distribution [60, 61], illustrated in Figure 5. For the Franka kitchen benchmark, we define the action space as the product of manifolds $\mathcal{M}=\mathbb{R}^{3}\times\mathcal{S}^{3}\times\mathbb{R}^{2}$, whose components represent the end-effector position, the end-effector orientation (encoded as quaternions), and the gripper fingers position. Regarding the real robot tasks, the action space is defined similarly as the product of manifolds $\mathcal{M}=\mathbb{R}^{3}\times\mathcal{S}^{3}\times\mathbb{R}^{1}$, whose components represent the position, orientation (encoded as quaternions), and opening of the gripper. For the Franka kitchen benchmark and the real robot tasks, the Euclidean and hypersphere parts employ Euclidean Gaussian distributions and a wrapped Gaussian distribution, respectively. Notice that this choice of prior distributions follows common practice in flow matching literature, where Gaussian and uniform priors are widely adopted due to their simplicity and effectiveness [16]. While more sophisticated approaches could be explored, for example using Gaussian Process as prior distributions [62], or learning task-dependent priors as in D-Flow [63], the choice depends on the problem at hand. We focus here on standard priors to maintain simplicity and consistency across tasks, while isolating their impact on our framework.
@@
State-based simulation tasks
@@
[…]
Vision-based simulation tasks
[…]
@@
For all experiments, we optimize the network parameters of RFMP and SRFMP using AdamW [64] with a learning rate of $\eta=1\times 10^{-4}$ and weight decay of $w_{d}=0.001$ based on an exponential moving averaging (EMA) framework on the weights [65] with a decay of $w_{\text{EMA}}=0.999$. We set the SRFMP parameters as $\lambda_{{\bm{x}}}=\lambda_{\tau}=2.5$. Table I summarizes the image resolution, number of parameters, and number of training epochs used in each experiment. Note that the policies predicts a sequence of actions $\bm{a}$ over a given time horizon $T_{p}$. Therefore, the total action space corresponds to the Cartesian product of manifolds computed over the prediction horizon defined for each task. Table II reports the total task dimensionality $\dim(\bm{a})$ as a function of the action space dimension $\dim(\bm{a}^{s})$ and the prediction horizon $T_{p}$ for all experiments considered in the paper. Notably, the Franka Kitchen benchmark involves learning a policy that predicts a $1944$-dimensional action sequence, enabling us to test the scalability of our method in high-dimensional, long-horizon settings. We use an action horizon $T_{a}=T_{p}/2$, and an observation horizon $T_{o}=2$ for all tasks.
@@ V-A2 Baselines
In [2], DP is trained using either Denoising Diffusion Probabilistic Model (DDPM) [12] or Denoising Diffusion Implicit Model (DDIM) [13]. In this paper, we prioritize faster inference and thus employ DDIM-based DP for all our experiments. We train DDIM with $100$ denoising steps. The prior distribution is a standard Gaussian distribution unless explicitly mentioned. During training we use the same noise scheduler as in [2], the optimizer AdamW with the same learning rate and weight decay as for RFMP and SRFMP. Note that DP does not handle data on Riemannian manifolds, and thus does not guarantee that the resulting trajectories lie on the manifold of interest for tasks with Riemannian action spaces, e.g., the Sphere Push-T, the Franka kitchen benchmark, and the real-world robot experiments. In these cases, we post-process the trajectories obtained during inference and project them on the manifold. In the case of the hypersphere manifold, the projection corresponds to a unit-norm normalization.
We also compare RFMP and SRFMP against CP [3] on the Robomimic tasks with vision-based observations. For a fair comparison, we retrained CP from scratch using the code provided in [3]. We pretrain the DP teacher policy for $100$ epochs and subsequently distilled the student policy for another $100$ epochs, following the CP training procedure [3]. Notice that this increases the total training time, effectively doubling the number of epochs, compared to RFMP, SRFMP, and DP. The neural network architecture used in DP and CP to model the diffusion noise is identical to the one used in RFMP and SRFMP. This isolates the influence of the learning algorithm, eliminating the architectural aspects as confounding factors.
TABLE II: Action space dimensionality for all experiments.
Task $T_{p}$ $\dim(\bm{a}^{s})$ $\dim(\bm{a})$
Euclidean Push-T $16$ $2$ $32$
Sphere Push-T $16$ $2$ $32$
Robomimic tasks $16$ $7$ $112$
Franka Kitchen $216$ $9$ $1944$
Real-world experiments $16$ $7$ $112$